### v2

# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

Abstract
Vision Language Models (VLMs) can be trained more efficiently if training sets can be reduced in size. Recent work has shown the benefits of masking text during VLM training using a variety of approaches: truncation, random masking, block masking and syntax masking. In this paper, we show that the best masking strategy changes over training epochs and that, given sufficient training epochs. We analyze existing text masking approaches including syntax masking, which is currently the state of the art, and identify the word frequency distribution as important in determining their success. Experiments on a large range of data sets demonstrate that syntax masking is outperformed by other approaches, given sufficient epochs, and that our proposed frequency-based approach, called Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) has numerous advantages. The benefits are particularly evident as the number of input tokens decreases.
Keywords Vision-Language Model  $\cdot$ Multimodal Data  $\cdot$ Word Frequency Masking.
1 Introduction
Vision-Language Models (VLMs) learn embeddings that represent both visual and textual modalities simultaneously, and are known for the ease with which their strong performance transfers from one domain to another [1, 2, 3, 4, 5, 6, 7, 8]. CLIP (Contrastive Language-Image Pre-training) [3, 9, 10] is a VLM that is trained on a very large number of image-text pairs collected online, where the text is considered a “natural language” annotation for the image.
However, pre-training CLIP models typically requires thousands of GPU-days [3, 9]. To accelerate the pre-training process, FLIP, A-CLIP, and CLIPA use masking strategies for image and text tokens when pre-training CLIP models [8, 11, 12]. In addition to the speed-up, masking makes it possible use large batch sizes, which enhances the benefits of contrastive learning and has been considered as a form of regularization, reducing the risk of overfitting [8]. For these reasons, despite the loss of some information from the original image and text, masking strategies have the potential to improve model performance.
Figure 1: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16. The models are pre-trained on CC12M for 45 epochs with 75% text masking ratio. We also use image masking (75%) to speed up pre-training. All approaches are subsequently fine-tuned with one epoch on the full, unmasked data.
Currently, the state of the art in text masking is represented by CLIPA [12], which compares four masking strategies: truncation, random word selection, random block selection, and syntax-based selection, i.e., masking on the basis of part-of-speech (POS) information. The conclusion is that syntax masking is the strategy that, overall, provides the best quality. In this paper, we carry out a set of experiments that demonstrate that this conclusion must be revisited. Specifically, as shown in Fig. 1, if enough epochs are used, other forms of masking substantially outperform syntax masking. The conclusion of CLIPA [12] misses the weakness of syntax masking because it trains with a limited number of epochs (6.4 epochs).
The special focus of our paper is a frequency masking approach that we introduce: CLIPF (CLIP Frequency masking). In Fig. 1, we see that with enough epochs CLIPF and CLIPA block masking perform similarly. We provide an analysis that demonstrates why CLIPF, block masking, and random masking emerge as clearly superior to syntax masking and truncation. The analysis studies the distribution of words by frequency of occurrence in the training data after masking. It shows that good masking strategies maintain helpful properties of this distribution and that syntax masking distorts the distribution of words over POS categories. These observation leads to the main insight of our paper, namely, when masking text for VLM pre-training, what you need for best results is to pay attention to the impact of masking on word frequency.
Our CLIPF approach is useful in practice because it performs relatively well in the range of 16-40 epochs. This range is a typical number of epochs to train a system in practice. For example, Open_CLIP [9] trains on 30 epochs. In Fig. 1, we see that in this range CLIPF outperforms other masking strategies. We carry out a set of experiments with 30 epochs to shed further light on the role of frequency in masking by comparing CLIPF vs. the different variants of CLIPA masking on zero-shot tasks and on image retrieval. We also investigate training on a larger data set.
The history of text masking in language model training dates to [13], where subsampling of frequent words in the training text was used to improve the performance of skip-gram language models. The study of text masking for VLMs got off to a slow start, since initially [8] reported that text masking, although it promotes efficiency, deteriorated performance. The CLIPA [12] work is important because it expanded the types of masking studied and demonstrated that masking can indeed improve VLM performance. To our knowledge, no previous work has attempted to discover the reason that certain forms of text masking outperform others or identified the importance of the way in which text masking approaches impact word frequency.
In sum, our paper makes the following contributions to the study of text masking in pre-training VLMs:
  • We point out the importance of the number of training epochs when comparing text masking strategies.
  • We show that word frequency distribution after masking differentiates between successful and less successful text masking strategies.
  • We carry out an in-depth comparison between existing masking strategies (CLIPA variants) and an approach that carries out masking by using frequency information directly, CLIPF. Our experiments show that CLIPF consistently outperforms syntax masking, previously considered the best choice for masking and provide a systematic view of how text masking can accelerate pre-training of VLMs while at the same time maintaining or improving VLM performance.
Our work is reproducible because it uses training data sets that are openly available and small enough to allow researchers with limited computational resources to repeat our experiments, and because we release our code [^1] https://github.com/MingliangLiang3/CLIPF.git.
2 Related work
2.1 Vision-Language Models
Vision-language models (VLMs), such as CLIP, ALIGN, LiT, and FLIP [3, 4, 10, 8], learn image embeddings from large-scale natural language supervision using contrastive learning [14], which is important for making models transferable. Because CLIP does not release its pre-training dataset, OpenCLIP pre-trained models on several large-scale public datasets [15, 16, 17, 18] to reproduce CLIP. This implementation adheres to the scaling law observed in Neural Language Models [19] and Vision-Language Models [3], where loss decreases according to a power-law relationship with model size, dataset size, and computational resources utilized during training. While large language models (LLMs) have advanced rapidly, we focus on vision-language models (VLMs) like CLIP, as they are essential tools for bridging visual and textual modalities in current research. Although CLIP and OpenCLIP are pre-trained on very large-scale datasets, recent works like DeCLIP [20] and SLIP [5] demonstrate that Vision-language Models (VLMs) can be trained efficiently with smaller datasets, enabling researchers with limited computational resources to contribute. Moreover, masking image patches or reducing image resolution during VLM pre-training also substantially decreases computational costs while maintaining or enhancing performance, as shown by techniques like FLIP, RECLIP, CLIPA [8, 21, 12].
2.2 Text masking for VLM training
Masking text presents another viable option to speed up the pre-training of VLMs, and is particularly effective alongside strategies that employ a high ratio of image patch masking. FLIP further explored text masking techniques, namely, random and truncation masking, to enhance efficiency [8].
While random text masking involves masking text arbitrarily, truncation masking in FLIP strategically eliminates padding tokens and retrains the words that preceded them.
Subsequently, CLIPA [12] introduced block masking and syntax masking. As already mentioned in Section 1, CLIPA [12] shows that text masking during training can improve the performance of a VLM. CLIPA syntax masking prioritizes the retention of nouns, followed by adjectives and then other words. In [12], syntax masking outperforms the truncation, random, and block approaches. However, [12] does not vary the number of training epochs. An important contrast with our work, as previously mentioned, is our attention to the number of training epochs. Further, CLIPF, as we will see, counters the danger of giving overly high priority to nouns, leading to overfitting, and also avoids overly suppressing words of the POS category “Other”.
An initial effort to exploit frequency information for text masking during VLM training was made by SW-CLIP [22]. Unfortunately, SW-CLIPF introduces a frequency cutoff threshold, which reduces performance. In order to exploit frequency information effectively each training text needs to fill as many available input tokens as possible, as is done in CLIPF.
3 Text Masking
In this section, we describe the text masking strategies that we study in this paper: the current state of the art, namely, the text masking strategies of CLIPA [12] (Section 3.1) and our frequency masking strategy CLIPF (Section 3.2). All these strategies accelerate pre-training by masking specific words in the text.
Consider a text $T_{i}$ consisting of words $\{w_{1},w_{2},w_{3},\ldots,w_{n}\}$, where $n$ represents the total number of words in $T_{i}$. Text masking involves selecting a subset of these words to create a new version of the text, denoted as
$T_{new\_i}=f(T_{i}\{w_{1},w_{2},w_{3},\ldots,w_{n}\}).$ (1)
This function $f$ determines which words are masked to get the modified text $T_{new\_i}$. And, we define the number of input tokens as $k$. It is important to note that the number of padding tokens remains consistent across all these strategies.
3.1 Existing Text Masking Strategies (CLIPA)
The four text masking strategies from CLIPA [12], which we also investigate here, are:
Random Masking: In the random text masking strategy, the function $f$ operates to randomly select words from the text $T_{i}$. Each word $w_{i}$ has an equal probability of being masked, independently of its position or frequency in the text.
Truncation Masking: In truncation masking, the function $f$ reduces the text by masking words from the end part of the text. The number of text tokens retained depends on the input length of the text encoder.
Block Masking: Block masking is similar to truncation masking but starts at a random index within the text. If the text tokens are longer than the input tokens, we choose a random starting index $j$ that is less than the length of the text tokens minus the length of the input tokens $k$.
Syntax Masking: To implement syntax masking [12], each word $w_{i}$ in the text $T_{i}$ is first tagged using a natural language processing tool like NLTK [23] to identify its part of speech. The words are then sorted according to their syntax role. This ordering strategy $O(w_{i})$ reflects the importance of nouns, adjectives, and verbs in the text. Words are masked based on their assigned order, with priority given to retaining nouns, followed by adjectives, verbs, and then other parts of speech. This method ensures that the core semantic information of the text is less likely to be masked, especially when learning the relationship between the image and the text, as nouns are more relevant to the object of the image.
3.2 Word Frequency Masking (CLIPF)
As mentioned in the introduction, subsampling of frequent words was introduced by [13] in the natural language processing (NLP) domain to accelerate the learning process for word representations. The goals are to speed up training and to address the imbalance between frequent and infrequent words in the training corpus. Our CLIPF method, uses Formula 2, which was introduced in [13], to determine the probability of masking a word during training. Thus, the complete formula to calculate the probability $P(w_{i})$ is:
$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (2)
Here, the masking probability of the word $w_{i}$ is $P(w_{i})$, $t$ is the threshold, and $f(w_{i})$ is the frequency of the word. This formula scales the word frequency $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones. It also preserves the overall ranking of word frequencies while aggressively masking high frequency words, controlled by the threshold $t$ [13].
4 Experimental Setup
Our implementation follows CLIP [3], OpenCLIP [9], FLIP [8] and CLIPA [12]. In this section, we present the details of our experimental setup.
Dataset We pre-train the models on Conceptual Captions 3M (CC3M) [24], Conceptual 12M (CC12M) [15]. These datasets were selected due to their diverse range of real-world concepts and scenes. Popular VLMs such as OpenCLIP, FLIP, DeCLIP, SLIP, A-CLIP and CLIPA [9, 8, 20, 5, 11, 12] also employ these datasets for training, leveraging their extensive and diverse content. Due to the expiration and inaccessibility of some URLs in the training set, we have downloaded 2.7 million image-text pairs for CC3M and 9.3 million image-text pairs from CC12M. The characteristics of the dataset are presented in Table 1. The CC12M dataset is 3.42 times larger than the CC3M dataset, with captions that are on average 2.15 times longer. Since the number of input tokens is fixed, the impact of different text masking ratios varies across datasets due to differences related to caption length. In an additional set of experiments, we provide results training on LAION-400M [17], a larger data set with about 300M image-text pairs.
Table 1: Dataset Statistics for Pre-training datasets. Caption length refers to the number of words in the text.
Dataset Samples Total words Caption length
CC3M 2721877 28022082 10.30 $\pm$ 4.72
CC12M 9295444 205716854 22.15 $\pm$ 17.20
Architecture For the image encoder, we use ViT-B/16 [25] architectures with global average pooling and learnable positional embeddings. The input image size is 224 for CC3M and CC12M. For the text encoder, we use a Transformer-based model [26] and byte-pair encoding with a 49K token vocabulary. We chose these encoders because they are the largest encoders that we can use given our resources. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
We trained the models using 8 RTX A5000 GPUs with the same settings to ensure consistent conditions across all models. We experiment with different text token input sizes, namely, 32, 16, 8, 6, and 4 text tokens. As the input size gets smaller, we can increase the batch size, maximizing computational memory usage. For CC3M and CC12M, the batch sizes corresponding to the input sizes are: 664, 832, 896, 928, and 944. Note that CLIP (and therefore CLIPA and CLIPF) is trained using contrastive learning, which benefits from larger batch sizes. Note also that although we used different text masking with the same settings and batch sizes, syntax masking was slower than the other text masking strategies when conducting the experiments because POS processing for each word is time-consuming.
We pre-train the model using the InfoNCE loss [14] with a learnable temperature parameter $\tau$ [27, 3]. To classify images, we calculate the cosine similarity between the image and text embeddings.
Table 2: Details of the pre-training and fine-tuning setups.
Config Pre-training Fine-tuning
optimizer AdamW [28] AdamW [28]
learning rate 1e-3 1e-5
weight decay 0.2 0.2
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.98$ [29] $\beta_{1},\beta_{2}=0.9,0.98$ [29]
learning rate schedule cosine decay [30] cosine decay [30]
warmup steps 10k 10%
epoch 30 1
$t$ $10^{-6}$
$\tau$ 0.07
numerical precision amp amp
RandomResizedCrop (50, 100) (50, 100)
Table 3: The number of image and text tokens that are processed during pre-training (CC3M and CC12M).
Image masking Text masking Image tokens Text tokens Total Percentage
0% 0.00% 196 32 228 100%
75% 0.00% 49 32 81 35.53%
50.00% 49 16 65 28.51%
75.00% 49 8 57 25.00%
81.25% 49 6 55 24.12%
87.50% 49 4 53 23.25%
Training Following CLIP, OpenCLIP [3, 9], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. Details of the pre-training configuration are given in Table 2.
Similar to FLIP, to speed up training, we apply 75% image masking to the image encoder as the baseline model. As a result, the speedup is about 4$\times$ compared to training without image masking, while the reduction in the zero-shot performance of ImageNet-1K classification remains within reasonable bounds.
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model. As shown in Table 3, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking. Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff. The details of the fine-tuning configuration are provided in Table 2.
Table 4: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
Models Masking Image Tokens Text Tokens CC3M CC12M
pre-train fine-tune pre-train fine-tune
CLIP 197 32 18.6 36.6
FLIP 49 32 14.1 14.2 32.0 33.7
CLIPA truncation 49 16 13.8 13.8 32.8 32.8
random 13.9 13.9 33.7 34.3
block 13.9 13.9 34.3 34.8
syntax 13.3 12.8 32.2 34.4
CLIPF frequency 14.0 14.0 35.5 36.0
CLIPA truncation 49 8 10.8 12.0 25.4 28.4
random 17.6 17.4 34.5 36.9
block 16.2 16.6 35.5 37.9
syntax 17.2 17.5 28.5 35.0
CLIPF frequency 16.8 17.0 36.6 39.3
CLIPA truncation 49 6 8.4 9.4 15.3 23.2
random 12.8 17.9 26.9 34.6
block 12.9 17.0 28.6 35.9
syntax 12.2 15.7 25.2 32.6
CLIPF frequency 14.4 18.2 30.3 37.8
CLIPA truncation 49 4 3.8 8.2 5.3 19.8
random 5.4 14.6 14.0 27.1
block 7.5 14.5 18.7 26.6
syntax 8.9 13.0 14.2 24.6
CLIPF frequency 10.9 16.0 17.0 30.9
CLIPF frequency 98 8 19.6 18.7 39.8 41.0
Evaluation setup Following CLIP [3], FLIP [8], and CLIPA [12], several classification benchmarks were used. Among these benchmarks is ImageNet-1K, a widely recognized dataset in computer vision. It is frequently used for image classification and VLM tasks and comprises 50K validation samples across 1K different classes. We filled each class into the templates provided by CLIP [3] to calculate the average of the text embeddings. We use the same evaluation settings as CLIP [3] to evaluate the other downstream datasets.
5 The Importance of Frequency Distributions
Our work is motivated by the patterns that we observe when VLMs are trained for additional epochs, shown in Fig. 1. This experiment motivates revisiting CLIPA [12], which was trained only on 6.4 epochs, in two respects. First, as already mentioned in the introduction, the conclusion that syntax masking is the best masking approach does not hold. Second, truncation masking, which looks relatively promising in low numbers of epochs, does not perform well given more epochs. In this section, we analyze the impact of masking on the training data in order to discover the reason behind this behavior.
Figure 2: Comparison of word masking probabilities for various methods. In this example, we keep four words for truncation and syntax. Numbers indicate the probability of a word being masked. Truncation keeps the first 4 words; Random and Block mask each word with a 50% probability; Syntax prioritizes retaining nouns, followed by adjectives, then others; Frequency masks words based on the frequency of the words.
We start with a basic illustration of the differences between task-masking strategies is provided in Figure 2. Truncation consistently retains the beginning of the text. With the random (random word selection) and block (random block selection) methods, each word has an equal probability of being masked, with block masking maintaining word neighborhoods. Syntax masking prioritizes the retention of nouns, followed by other part-of-speech (POS) categories.
Figure 3: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies. We set the text length after text masking to 6. The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
Next, we look at how the different masking strategies impact the frequency distribution of words in the training data. Fig. 3 shows the distribution of words in the text before and after applying all of the text masking strategies. We see that in general, there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency. Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed. Most masking approaches preserve the general shape of this distribution: The frequency of a word in the masked training data tracks its frequency in the original data. The exception is syntax masking, which shows peaks and pits for specific words. Specifically, because syntax masking retains more nouns, the articles “the” and “a” are less frequent, but the nouns create spikes. The rationale for prioritizing nouns is solid. Nouns are closely related to the depicted content of images and, for this reason, are considered to be more helpful. Indeed, the helpfulness of nouns for the VLM is reflected in the fact that syntax masking learns faster than other text masking strategies. In Fig. 1, we saw that syntax masking outperforms the other masking approaches during the early epochs. However, retaining too many nouns can have two adverse effects. First, the model can overfit to certain high frequency nouns. In Fig. 3, we see that CLIPA syntax masking retains more occurrences of “person,” “illustration,” and “stock” than the other strategies. Second, retaining nouns means necessarily dropping words from other parts-of-speech (POS) categories, which may be important, as discussed in more detail below.
Considering the CLIPA truncation masking strategy, we see in Fig. 1 that the word frequency distribution of the truncated training data shows a difference with the other masking strategies. Specifically, truncation is the top line on the left of the graph, which means truncation is making the most highly frequent words in the training data relatively even more frequent. The fact that truncation impacts word frequency might seem surprising at first, but it can be explained if we consider that the probability of a word occurring changes with sentence position. With truncation, typical text beginnings, e.g., “A person is…”. remain over-represented in the data.
To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 5. Consulting the column “Other”, we see an interesting pattern. Syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but truncation, random and block do not. CLIPF appears to provide a happy medium. CLIPF retains more nouns than CLIPA truncation, random, and block, but not so many that the “Other” category becomes poorly represented.
Table 5: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking. The dataset is CC12M, and we retain 6 words for each text after applying text masking.
Masking NN (%) JJ (%) VB (%) OTHER (%)
Before masking 50.30% 4.98% 5.18% 39.55%
Truncation 51.72% 5.17% 5.91% 37.20%
Random 51.21% 4.82% 5.04% 38.93%
Block 49.37% 4.74% 5.24% 40.66%
Syntax 88.53% 2.69% 2.92% 5.87%
CLIPF 60.43% 5.06% 6.20% 28.24%
The importance of “Other” words is surprising because the category includes many words, which like articles, lack semantic content and are sometimes considered to have no direct connection to images. In fact, in many NLP applications, articles and other “stop words” are simply removed. However, since the CLIP text encoder is a transformer, technically it should be able to exploit the context of content words, particularly words in “Other” category. Note that with syntax masking a phrase like “smoking in the room” in the training data would always be reduced to “smoking room”, which has a different meaning. However, with the other masking approaches, “in the” would not be consistently deleted. Note specifically that block masking retains entire phrases, without always using the beginning of the sentence as truncation masking does. Also, frequency masking (CLIPF) retains entire phrases in some cases, more than would be expected with random masking.
In sum, when trained with enough epochs, overfitting effects with syntax masking occur. Our analysis has related this problem to peaks in the word frequency distribution caused by over-retention of nouns. Also, with enough epochs, truncation masking overfitting effects occur, which our analysis has related to distortion of the word frequency distribution caused by oversampling of the beginning of sentences. In sum, our analysis of word frequency distributions in the training data provides insight into why the performance of syntax and block masking falls off in Fig. 1. Further, if word frequency is indeed important, then masking that controls word-frequency directly, such as our CLIPF, can be considered to be a sensible and transparent approach.
6 Experimental Results
In this section, we carry out a systematic investigation of the existing text masking approaches from CLIPA [12] and our CLIPF approach. Our results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF. Our experiments use 30 epochs, following Open_CLIP [9]. We note that VLM training work has shown performance improvements up to 100 epochs [5, 11] and we would expect that with more epochs the performance would also improve. However, the improvement beyond 30 is expected to be small, predictable and not necessarily important in an operational setting where conserving compute is of concern.
6.1 Zero-shot Performance on ImageNet-1k
Table 4 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine-tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores. In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse. Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance.
It is particularly interesting to observe from the results in Table 4 the fact that models trained with text masking (CLIPA and CLIPF) are competitive with, or superior to, models trained on the original, unmasked text data (CLIP, FLIP). Note that at a text masking ratio of 75% (corresponding to 8 tokens), the masking approaches are 1.3 $\times$ faster than FLIP and 4.0 $\times$ faster than CLIP.
We first consider the comparison with CLIP. On CC3M, CLIPF approaches the performance of the original CLIP approach with a 81% text masking ratio (4 input tokens). On CC12M, CLIPF outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens). Fine-tuned CLIPF with a 75% masking ratio achieves an accuracy of 39.3% compared to 36.6% for the original CLIP trained on the full data.
Next, we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens). On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios. On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens). Additionally, as shown in the last row of Table 4, if we double the number of image tokens, then CLIPF substantially beats the original CLIP (41.0% vs 36.6%). However, it also notes that more resources are required (Tokens processed from 25% to 46%).
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. Recall that the reduction of the number of input tokens allows for an increase in the batch size. As we reduce the number of tokens from 8 to 6 to 4, syntax masking and truncation do not perform as well as the other text masking strategies. We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
6.2 Zero-shot Robustness Evaluation
Table 6: Zero-shot robustness evaluation. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens ViT-B/16
IN-A IN-O IN-R IN-S IN-V2 ON
CLIP 197 32 8.97 37.85 49.11 25.70 31.48 24.20
FLIP 49 32 6.90 40.30 36.90 17.60 28.70 18.10
CLIPA truncation 49 16 7.10 37.70 36.40 17.50 27.40 17.20
random 8.05 38.60 40.00 20.34 29.61 20.84
block 8.05 39.60 39.17 19.31 29.70 19.72
syntax 8.00 40.70 39.41 18.61 29.17 19.96
CLIPF frequency 8.28 39.10 40.86 20.48 30.78 21.15
CLIPA truncation 49 8 6.71 34.05 37.18 15.99 24.49 15.98
random 9.16 38.15 43.35 22.53 31.24 22.17
block 9.05 39.95 43.32 22.58 32.42 21.58
syntax 8.50 38.40 39.90 19.50 29.80 19.50
CLIPF frequency 9.56 39.90 45.76 23.62 34.16 21.59
CLIPA truncation 49 6 6.70 26.60 33.20 13.30 20.20 14.80
random 8.10 36.10 41.80 21.10 30.20 19.80
block 8.32 38.20 42.99 21.73 30.82 19.76
syntax 7.40 34.10 39.60 18.50 28.60 19.80
CLIPF frequency 8.69 38.55 45.09 23.45 32.59 20.04
CLIPA truncation 49 4 5.30 21.70 25.40 9.00 17.10 11.20
random 6.10 30.30 33.80 14.90 23.50 15.30
block 6.08 30.55 34.69 15.54 22.99 15.10
syntax 6.50 29.10 32.40 13.30 22.00 16.00
CLIPF frequency 7.05 32.40 39.90 18.94 26.72 17.14
Following the evaluation of CLIP and FLIP [3, 8], we test the masking strategies across data sets. Table 6 presents results for zero-shot classification for six commonly used data sets. The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [31] by 1.28%, ImageNet-O [32] by 0.80%, ImageNet-R[33] by 4.9% , ImageNet-Sketch [34] by 3.14%, ImageNet-V2 [35] by 3.38%, and ObjectNet [36] by 0.44%. Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases. However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in 6.1.
Table 7: Zero-shot Image-Text Retrieval. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Text Tokens Text Retrieval Image Retrieval
Flickr30k COCO Flickr30k COCO
R@1 R@1 R@1 R@1
CLIP 32 62.62 35.54 45.42 24.22
FLIP 32 54.73 29.34 38.62 20.88
CLIPA truncation 16 44.67 25.54 34.99 19.64
random 16 58.48 32.36 43.61 22.81
block 16 56.51 30.82 44.06 22.66
syntax 16 54.54 29.60 41.16 21.40
CLIPF frequency 16 57.89 31.52 42.72 22.57
CLIPA truncation 8 30.47 16.92 23.96 12.78
random 8 58.58 36.24 43.79 23.16
block 8 60.06 35.88 45.98 24.65
syntax 8 50.30 29.64 38.46 20.28
CLIPF frequency 8 58.68 34.74 44.58 23.31
CLIPA truncation 6 32.25 16.18 22.39 11.93
random 6 55.23 32.58 42.47 21.34
block 6 54.44 33.12 41.68 21.73
syntax 6 46.15 26.78 34.08 17.99
CLIPF frequency 6 56.02 32.32 41.05 21.28
CLIPA truncation 4 30.97 17.44 21.09 11.07
random 4 41.62 24.14 30.26 15.63
block 4 40.83 23.56 30.77 14.98
syntax 4 38.66 21.68 26.08 14.14
CLIPF frequency 4 45.07 25.60 31.72 15.96
6.3 Zero-shot Image-Text Retrieval
Table 7 shows image/text retrieval results on Flickr30k [37] and COCO [38]. In general, we see that text masking approaches are able to improve over FLIP. When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text. Here, we do not see consistent improvement over CLIP, although some cases come lose. Note again that we are testing here only with 49 image tokens and 98 would fare better, with additional computational cost.
In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies. The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets. The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used. Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
We point out that for Image-Text Retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking. Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval. This gap is wider than what was observed with Zero-shot Classification. We conjecture that the bad performance is due to the Image-Text Retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects. For this reason, the fact that syntax masking retains a high number of nouns, as shown in Section 5, is particularly harmful.
Table 8: Zero-shot accuracy on more classification datasets. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Method
Text Tokens
Food-101
CIFAR-10
CIFAR-100
CUB200
SUN397
Cars
Aircraft
DTD
OxfordPets
Caltech-101
Kinetics700
Flowers102
MNIST
STL10
EuroSAT
Resisc45
GTSRB
KITTI
Country211
PCAM
UCF101
CLEVR
HatefulMemes
SST2
ImageNet
CLIP 32 45.46 66.42 32.45 7.90 49.27 16.17 2.28 15.48 57.51 71.91 24.84 2.12 8.77 91.29 19.08 37.38 5.09 29.81 4.45 50.05 40.52 21.84 52.66 48.76 36.61
FLIP 32 36.26 58.09 26.89 8.73 38.56 8.16 2.98 19.95 51.85 67.17 21.57 2.94 9.35 81.89 18.02 32.30 9.68 21.39 3.63 50.26 36.51 14.67 51.37 50.08 33.70
CLIPA truncation 16 35.53 61.24 26.54 9.42 39.98 7.28 2.57 15.80 46.85 68.28 20.47 1.94 8.32 81.89 18.48 33.75 2.33 17.98 3.79 58.56 30.85 13.31 53.00 50.08 32.83
random 16 38.48 53.61 29.46 7.97 41.92 8.53 2.79 21.54 56.71 69.67 22.88 2.23 9.73 84.86 17.90 36.41 7.28 32.95 4.77 55.00 35.29 16.38 49.93 50.08 34.26
block 16 38.69 57.51 29.71 8.23 44.54 9.85 2.45 21.33 54.01 70.61 23.00 2.79 10.09 85.04 31.88 29.98 9.07 25.07 4.59 60.68 38.07 18.72 51.24 49.64 34.78
syntax 16 35.94 58.78 31.86 8.91 41.92 8.39 1.53 17.07 47.05 67.59 21.34 3.10 10.45 76.92 26.68 32.65 8.73 39.37 4.23 49.96 32.91 15.55 53.30 49.92 34.41
CLIPF frequency 16 39.47 64.43 32.55 10.15 45.51 9.22 2.43 21.70 55.62 70.69 23.08 2.77 9.56 83.00 21.02 33.30 7.20 27.47 4.12 50.24 37.72 13.70 52.70 49.92 35.96
CLIPA truncation 8 36.28 61.99 32.24 8.65 44.40 5.56 2.04 18.56 45.67 65.93 22.68 2.52 10.51 87.96 34.80 29.37 3.35 46.06 4.59 51.55 32.38 12.41 53.46 50.08 30.30
random 8 45.10 68.54 34.04 9.03 51.54 8.64 3.23 30.53 59.29 70.06 27.29 4.21 9.82 90.46 28.72 32.54 5.99 30.61 7.95 58.00 42.90 12.87 53.40 50.08 37.25
block 8 43.70 66.66 32.62 8.08 51.07 9.27 2.54 28.99 57.88 70.89 26.72 2.65 18.10 90.84 31.88 36.94 9.02 33.36 7.52 61.91 40.10 14.66 51.80 53.27 37.88
syntax 8 37.92 65.25 31.72 8.28 45.74 8.73 2.96 17.07 46.54 68.73 23.52 2.59 11.30 88.30 29.90 30.81 6.47 30.82 5.02 54.62 36.58 12.51 52.74 50.08 34.97
CLIPF frequency 8 45.75 59.00 30.76 10.60 50.06 10.32 3.20 29.89 67.54 73.35 26.96 4.09 10.10 90.54 20.80 36.30 7.16 40.31 7.10 57.23 43.09 12.27 54.54 50.03 39.30
CLIPA truncation 6 23.54 60.27 23.52 5.70 36.54 3.82 2.03 10.32 31.12 61.15 17.77 1.86 9.94 85.25 20.62 23.68 7.13 16.71 3.45 50.23 24.53 12.61 53.62 50.08 23.28
random 6 38.45 71.57 33.89 7.73 51.81 6.67 2.72 29.47 59.92 68.65 25.89 2.42 12.28 89.73 26.90 33.17 8.89 29.75 7.23 56.87 38.62 12.63 51.96 50.08 34.62
block 6 39.87 66.60 35.31 7.58 53.23 7.49 2.89 27.87 56.74 69.40 25.97 2.95 10.06 89.30 27.90 32.40 7.31 33.76 6.73 63.10 39.20 12.65 55.41 52.83 35.89
syntax 6 34.18 63.46 29.83 7.54 43.54 7.92 1.85 17.23 44.97 68.37 22.75 1.99 11.37 89.19 17.56 26.13 3.84 33.16 4.98 55.25 35.40 12.30 53.44 50.80 32.57
CLIPF frequency 6 45.33 67.64 32.83 9.98 50.86 9.61 2.19 31.33 66.03 69.24 26.60 2.94 10.71 90.98 33.64 33.22 8.93 36.70 7.32 59.96 40.60 12.45 54.15 50.08 37.81
CLIPA truncation 4 20.64 51.67 22.10 4.19 37.16 3.11 1.18 15.16 29.36 53.99 16.17 1.04 10.10 86.34 17.16 15.81 7.48 30.15 3.49 61.43 24.95 12.63 50.76 50.08 19.80
random 4 28.82 61.90 28.98 5.35 46.39 4.03 1.42 25.16 43.26 62.87 22.03 2.10 12.24 88.54 23.82 24.46 8.07 39.30 5.59 50.62 33.52 12.65 51.86 50.08 27.13
block 4 28.08 60.63 27.10 5.09 45.61 3.82 2.43 23.24 36.63 64.15 21.80 2.29 8.41 87.04 24.44 23.95 7.55 43.05 5.51 55.80 32.62 9.49 53.55 50.08 26.66
syntax 4 25.59 54.13 27.46 4.73 42.26 6.26 2.66 15.27 32.52 58.93 19.58 2.20 10.42 90.11 26.02 20.62 5.19 37.50 4.32 51.12 26.82 12.93 52.77 50.08 24.61
CLIPF frequency 4 36.09 69.76 31.14 8.41 47.72 7.01 2.24 25.64 54.23 64.55 23.55 2.95 10.18 90.25 20.78 26.52 6.90 36.63 6.40 55.71 35.71 12.65 49.63 50.08 30.93
6.4 Zero-shot Classification on More Datasets
In Table 8, we present evaluation results on additional datasets. With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies. Aligned with ImageNet findings, reducing text tokens from 16 to 8 significantly boosts performance. Improvements include SUN397 [39] (+6.39%), DTD [40] (+9.31%), OxfordPets [41] (+4.64%), STL10 [42] (+9.31%), EuroSAT [43] (+6.60%), UCF101 [44] (+6.68%). The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, their improvements are more limited.
We point out that the Describable Textures Dataset (DTD)[40] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[41] also contains many class names that are not nouns. On these data sets the performance of syntax masking is particularly low. Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets. This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking.
6.5 Scaling to the larger Dataset.
Finally, we carry out an exploratory experiment to gain insight into different masking strategies with data sets of different sizes. We carry out the largest experiment that is possible given our computational capacity, which was training on the LAION-400M [17] data set, using 4 text tokens, a smaller image size (112x112) and 50% image masking. We trained for 16 epochs. Previous works have shown that both smaller image sizes and image masking are effective for accelerating CLIP pre-training [21, 8, 45]. Results are shown in Fig. 4 for both top-1 accuracy (left) and top-5 accuracy (right). We see from these graphs that using $<$ 10 epochs would not provide a complete picture on how the different masking approaches perform relative to each other. Overall, CLIPF, the frequency based approach dominates. However, syntax is performing better than it did on smaller data sets, possibly because the large amount of data is compensating for the overemphasis of syntax masking on nouns. It is important not to conclude on the basis of this experiment that syntax masking will continue to dominate block masking. Looking at the top-5 performance, we see that these two approaches near each other already in the 16 epochs we are able to compute. In sum, this experiment shows that CLIPF is a good performer as data scales up, but that the number of epochs is important in understanding the strength of different masking strategies.
Figure 4: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies when pre-training on LAION-400M datasets. We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.
7 Conclusion and Outlook
This paper has provided a systematic overview of text masking approaches for VLM training, which accelerate training and have the potential to also improve performance. In this paper, we demonstrate that training epochs have a substantial impact on the performance of different text masking strategies. We conclude, contrary to the existing state-of-the-art, CLIPA [12], that syntax masking is not the best performing text masking approach and that truncation masking can be avoided. We carry out an analysis of word frequency distributions that shows that frequency is what you need to think about when text masking the training data of VLMs. Specifically, both syntax and truncation have visible effects on the word frequency distribution, making clear why they do not perform well.
We have introduced CLIPF, an approach for text masking based on word frequency that improves the performance of CLIP. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process. CLIPF is also interesting because it controls frequency directly and in doing so seems to balance nouns with words in the “Other” POS categories, in contrast to other approaches.
We carry out an extensive comparison among existing text matching strategies (CLIPA variants) and CLIPF that revals the strong performance of CLIPF in a variety of zero-shot classification tasks and when scaling to larger data sets. CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and other text masking strategies. It also delivers strong performance in under highly aggressive masking (i.e., using only 4 tokens per training text). Moving forward, the insights gained from this paper about the importance of frequency can be used to further optimize CLIPF to directly optimize the word distribution allowing for more tuning control when trading off VLM performance and VLM training speech during text masking.
Appendix A Text Masking Analysis
A.1 Text Masking Cases
As illustrated in Fig. 5, there is a clear relationship between word masking probability and word frequency. Frequent words have a higher masking probability compared to infrequent ones. Additionally, a smaller threshold $t$ leads to a smaller difference between the masking probabilities of frequent and infrequent words. Therefore, it was necessary to choose a relatively larger threshold to ensure that both frequent and infrequent words are effectively masked.
Figure 5: The curve of Formula 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of formular 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$,$10^{-8}$.
Table  9 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in Fig. 3, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content.
Table 10 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
Table 9: Example from the CC12M dataset illustrating the effect of various text masking strategies, reducing text length to 6 words. The original caption is in the first row, followed by masked variants. Parameter $t$ in Equation 2 is set to $10^{-6}$.
Method Resulting Text
Original Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
Truncation walk of the happy young couple
Random the happy and the is the
Block couple and siberian dog . the
Syntax walk dog man smiling head girl
CLIPF siberian handsome man hugging smiling red
Table 10: The probability of masking words in the text is calculated using the formula provided in Equation 2. The example is selected from CC12M [15]. The value of $t$ of formular 2 is set to $10^{-6}$.
Word Probability
walk 0.926171
of 0.992838
the 0.995064
happy 0.951695
young 0.957311
couple 0.941174
and 0.991695
siberian 0.75092
dog 0.960531
. 0.993678
A.2 Word Category Distribution Across Different Text Masking Strategies
As shown in Table 11, we provide the word type counts of Table 5. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
Table 11: Distribution of syntax counts and percentages before and after applying text masking. The dataset is CC12M and we retain 6 words for each text after applying text masking.
Masking NN Count JJ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
Before masking 103,469,117 10,245,828 10,649,943 81,351,966 50.30% 4.98% 5.18% 39.55% 205,716,854
Truncation 28,628,129 2,859,462 3,272,401 20,585,364 51.72% 5.17% 5.91% 37.20% 55,345,356
Random 28,334,735 2,666,847 2,790,600 21,550,517 51.21% 4.82% 5.04% 38.93% 55,342,699
Block 27,314,723 2,624,594 2,897,434 22,510,031 49.37% 4.74% 5.24% 40.66% 55,346,782
Syntax 48,989,317 1,483,666 1,618,088 3,245,892 88.53% 2.69% 2.92% 5.87% 55,336,963
SW-CLIP 34,645,367 3,346,676 4,073,825 6,062,345 71.98% 6.95% 8.47% 12.60% 48,128,213
CLIPF 33,516,439 2,803,409 3,473,119 15,666,310 60.43% 5.06% 6.20% 28.24% 55,459,277
Figure  6 shows the word cloud before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying frequency word masking, the word cloud in Table 6(b) shows a more balanced distribution of word frequencies compared to before the masking. The word “the" appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text. As depicted in Table 6(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
Fig. 6 shows word clouds before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying word frequency masking, the word cloud in Fig. 6(b) displays a more balanced distribution of word frequencies compared to before masking. The word “the” appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text. As depicted in Fig. 6(f), syntax masking tends to retain more nouns such as “person”, “stock”, and “illustration”, resulting in a different pattern from the others.
Refer to caption Refer to caption
(a) Before masking (b) Frequency
Refer to caption Refer to caption
(c) Truncation (d) Random
Refer to caption Refer to caption
(e) Block (f) Syntax
Figure 6: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words. The font size of the word represents the relative word frequency for each figure. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.3 Word Distribution Analysis for CLIPF
In addition, we analyzed frequency text masking strategies that varied according to the number of words used, as shown in Fig. 7. As the number of words decreased from 16 to 8 and then to 6, more frequent words were masked. However, reducing the number to 4 words led to a smaller vocabulary size, resulting in the loss of some important information. Consequently, the performance of the model pre-trained with 4 words was substantially lower compared to the model pre-trained with 6 words. We recommend setting the number of text words in a frequency-based text masking strategy to strike a balance between frequent and infrequent words and to maintain a larger vocabulary. Based on our experiments, the optimal number of text masking was found to be approximately 40-60% of the average length of the original text. This configuration helps achieve a balanced word distribution, which is beneficial for pre-training VLMs.
Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.4 Word Distribution of CLIPF and SW-CLIP
We also compare the word distributions of CLIPF and SW-CLIP after applying text masking. As illustrated in Fig. 8, SW-CLIP masks more frequent words based on its threshold, preventing it from fully utilizing the input slots of the text encoder. Consequently, SW-CLIP exhibits a flatter word distribution compared to CLIPF.
Figure 8: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
Appendix B More Results
More detailed results are presented in this section, including the results and learn curve of image-text retrieval.
B.1 The details of Image-text Retrieval
In Table 12, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
Table 12: Zero-shot Image-Text Retrieval, we evaluated CLIP, FLIP, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Models Masking Text Tokens Text Retrieval Image Retrieval
Flickr30k COCO Flickr30k COCO
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
FLIP 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
CLIPA truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
random 16 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
block 16 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
syntax 16 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
CLIPF frequency 16 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
CLIPA truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
random 8 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
block 8 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
syntax 8 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
CLIPF frequency 8 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
CLIPA truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
random 6 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
block 6 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
syntax 6 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
CLIPF frequency 6 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
CLIPA truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
random 4 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
block 4 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
syntax 4 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
CLIPF frequency 4 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
B.2 Learn Curved for Image-text Retrieval
We also investigate the zero-shot image-text retrieval performance over training epochs on the COCO and Flickr30k datasets. It also shows that the number of training epochs has a substantial impact on the performance, as illustrated in Fig. 9. And the performance gap between syntax masking and the other text masking on the zero-shot image-text retrieval task is larger than zero-shot classification task. This is because the zero-shot image-text retrieval task requires the model not only to understand the objects/nouns, but also to understand the logic of the text. Moreover, we present the zero-shot image-text retrieval performance across training epochs for models pre-trained on the LAION-400M dataset. As illustrated in Fig. 10, the performance patterns of models trained on LAION-400M are similar to those of previous models that pre-trained on the CC12M dataset. Specifically, CLIPF consistently outperforms syntax masking. Block masking achieves comparable performance to syntax masking and even surpasses it on the image-to-text retrieval task using the COCO dataset at epoch 16th.
Figure 9: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model for an additional epoch without image and text masking.
Figure 10: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on LAION-400M for 16 epochs with 25 image tokens and 4 text tokens to speed up training and fine-tune the model for 0.4 epoch with 224x224 image size and 32 text tokens.
Additionally, Fig. 11 illustrates the validation loss across pre-training epochs. It is evident that the validation loss for truncation and syntax masking methods begins to increase earlier compared to other text masking methods, suggesting that these methods may overfit to the data sooner.
Figure 11: The validation loss curve of COCO and Flick30k datasets over the pre-training epochs.
B.3 Learn Curved for DTD and OxfordPets Datasets
The learning curves for the DTD [40] and OxfordPets [41] datasets are presented because not all class names in these datasets are nouns. As shown in Table 13, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in Fig. 12. Moreover, the performance of syntax masking is very poor and almost close to truncation masking. OxfordPets is a dataset containing incomplete noun class names. In the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking. However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
Figure 12: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Figure 13: Zero-shot classification accuracy on DTD dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Figure 14: Zero-shot classification accuracy on OxfordPets dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
B.4 Learn Curved before fine-tuning
In Fig. 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning in Fig. 15 (30 epochs) and Fig. 16 (45 epochs). They have a similar pattern to the models before fine-tuning for every text masking strategy.
Figure 15: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training.
Figure 16: Zero-shot classification accuracy on ImageNet-1K over the 45 epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 45 epochs with image and text masking (75%) to speed up training.
B.5 Fine-tuning the models with More Epochs
The purpose of fine-tuning the models without image and text masking is to reduce the distribution gap between pre-training and evaluation. As demonstrated in FLIP and CLIPA, a small number of fine-tuning steps is usually sufficient to bridge this gap [8, 12]. We conducted additional experiments using two fine-tuning epochs, shown in Tab 13, to demonstrate that increasing the number of fine-tuning epochs does not substantially improve performance but instead leads to higher computational costs.
Table 13: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC12M for 30 epochs with 75% image masking for speedup (pre-train). Then, we fine-tune the model with one epoch (fine-tune-1) and two epochs (fine-tune-2) without image and text masking.
Models Masking Image Tokens Text Tokens CC12M
pre-train fine-tune-1 fine-tune-2
CLIP 197 32 36.6
FLIP 49 32 32.0 33.7 34.1
CLIPA truncation 49 8 25.4 28.4 28.5
random 34.5 36.9 37.1
block 35.5 37.9 38.0
syntax 28.5 35.0 35.5
CLIPF frequency 36.6 39.3 39.4
B.6 Comparison with SW-CLIP
In this section, we provide a comparison of CLIPF with SW-CLIP [22], which also uses frequency-based sampling but imposes a threshold on the frequency score. The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold. In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots. For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [22]. It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer. For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking. Subsequently, we fine-tuned both models with an additional epoch without any image or text masking. The results in Table 14 show that after fine-tuning CLIPF outperforms SW-CLIP.
Table 14: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Datasets Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SW-CLIP CC3M 49 6 14.9 16.8
CLIPF 14.4 18.2
SW-CLIP CC12M 49 8 35.9 38.5
CLIPF 36.6 39.3
B.7 Applying text masking on SigLIP
We also applied the text masking strategies to SigLIP; the results are presented in Tab. 15. CLIPF still achieves superior performance compared to other text masking strategies and without text masking. This further supports our conclusion that frequency-based masking is the more effective strategy.
Table 15: We apply text masking (75%) on SigLIP to pre-train on CC12M for 30 epochs with 75% image masking for speedup pre-training. Reporting the zero-shot classification accuracy on ImageNet-1K.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SigLIP 197 32 39.3
SigLIP 49 32 28.4 29.6
SigLIP truncation 49 8 20.4 27.1
random 29.7 31.7
block 30.5 32.7
syntax 25.9 31.6
SigLIP frequency 32.4 34.8
Appendix C Ablation
Table 16: We pre-train CLIPF on CC12M dataset across different thresholds in Equation 2. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
model Image tokens Text tokens Threshold ViT-B/16
pre-train fine-tune
CLIPF 49 8 1e-5 35.6 38.4
1e-6 36.6 39.3
1e-7 36.1 38.6
C.1 Threshold Analysis
To investigate the impact of the threshold $t$ in Equation 2 on model performance, we pre-trained models with varying thresholds. As shown in Table 16, thresholds of $51e-$ 5, $61e-$ 6, and $71e-$ 7 yield comparable results, all outperforming the other text masking strategies. This indicates that CLIPF is relatively insensitive to small variations in thresholds, which are intended to maintain the word masking probability within the range of 0 to 1. However, using too small a threshold reduces the differences in text masking probabilities; therefore, we do not recommend using an excessively low threshold.
Table 17: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
CLIPF frequency-token 49 8 36.0 38.0
frequency-word 49 8 36.6 39.3
C.2 Word and Token Analysis
Since Open_CLIP encodes text using byte-pair encoding (BPE) [9], some words are represented by multiple tokens. Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words. Table 17 compares the performance of both approaches. Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.